1. What is the partial pressure of nitrogen in a container that contains 5.82 mol of oxygen, 7.90 mol of nitrogen, and 5.85 mol of carbon dioxide when the total pressure is 621 mmHg?

A)

185 mmHg

B)

186 mmHg

C)

251 mmHg

D)

420 mmHg

E)

621 mmHg

2. Using the van der Waals equation, determine the pressure of 335.0 g of SO2(g) in a 5.70-L vessel at 661 K. For SO2(g), a = 6.865 L2 • atm/mol2 and b = 0.05679 L/mol. (R = 0.0821 L • atm/(K • mol))

A)

9.35 atm

B)

49.8 atm

C)

9.44 atm

D)

46.7 atm

E)

58.3 atm
Answer #1

1. 1. What is the partial pressure of nitrogen in a container that contains 5.82 mol of oxygen, 7.90 mol of nitrogen, and 5.85 mol of carbon dioxide when the total pressure is 621 mmHg?

Answer

The partial pressure is directly proportional to the mole fraction of the gas in the mixture of gas

Total moles of gas present = 5.82 + 7.9 + 5.85 = 19.57 moles

mole fraction of nitrogen = 7.9 / 19.57 = 0.403

so partial pressure of nitrogen = total pressure exerted by the gases X mole fraction of nitrogen

     = 621 mmHg X 0.403 = 250.263 almost equal to 251 mm Hg ( option c)